# UniAttn: Reducing Inference Costs via Softmax Unification for Post-Training LLMs

### v1

@@ Appendix G More Insights on Linear Compensation
$\mathbb{E}\bigl{[}\min_{X}||AX-B||\bigr{]}=\begin{cases}\sqrt{n(m-n)},&\text{if }m\geq n,\\ 0,&\text{if }m<n.\end{cases}$ (27)
@@
We give the proof for Theorem G.1 in the following subsections.
G.1 Preliminaries
To prove Theorem G.1, we adopt the notations in Theorem G.1 and prove a series of lemmas first.
Lemma G.2.
Let $P=AA^{+}$, $P$ is an orthogonal projection and $\text{Im}(P)=\text{Col}(A)$.
@@ G.1 Preliminaries
Lemma G.3.
$I-P$ is an orthogonal projection and $\text{Im}(I-P)=\text{Col}(A)^{\perp}$.
@@
First, we prove that $I-P$ is an orthogonal projection:
@@
$\langle\mathbf{v},\mathbf{d}\rangle=\langle(I-P)\mathbf{c},A\mathbf{w}\rangle=\langle\mathbf{c},(I-P)A\mathbf{w}\rangle=\langle\mathbf{c},(A-AA^{+}A)\mathbf{w}\rangle=\langle\mathbf{c},0\rangle=0$ (33)
@@
Conversely, suppose $\mathbf{u}\in\text{Col}(A)^{\perp}$. Recall Lemma G.2 that $P$ is an orthogonal projection and $\text{Im}(P)=\text{Col}(A)$, we have $P\mathbf{u}=0$. Therefore,
@@
From Lemma G.3, we can immediately conclude that:
Corollary G.4.
@@
$\text{rank}(I-P)=\dim\text{Im}(I-P)=\dim\text{Col}(A)^{\perp}=m-\text{rank}(A)$. Recall Theorem G.1 that every element from $A$ is sampled from a Gaussian distribution, so that by probability of 1 we have $\text{rank}(A)=\min(m,n)$, which leads to the conclusion of this corollary. ∎
@@
$\mathbb{E}\bigl{[}||Q\mathbf{b}||^{2}\bigr{]}=\mathbb{E}\bigl{[}\mathbf{b}^{T}Q^{T}Q\mathbf{b}\bigr{]}=\mathbb{E}\bigl{[}\mathbf{b}^{T}Q\mathbf{b}\bigr{]}=\mathbb{E}\bigl{[}\text{tr}(\mathbf{b}^{T}Q\mathbf{b})\bigr{]}=\mathbb{E}\bigl{[}\text{tr}(Q\mathbf{b}\mathbf{b}^{T})\bigr{]}=\text{tr}(Q\mathbb{E}\bigl{[}\mathbf{b}\mathbf{b}^{T}\bigr{]})$. As $\mathbf{b}\sim N(\mathbf{0},I_{m\times m})$, $\mathbb{E}\bigl{[}\mathbf{b}\mathbf{b}^{T}\bigr{]}=I$. Therefore, using the trace property of projectors:
@@
In turn, we can conclude that $\mathbb{E}\bigl{[}||Q\mathbf{b}||\bigr{]}=\sqrt{k}$. ∎
G.2 Proof
Finally, we give the proof for Theorem G.1.
@@ G.2 Proof
According to Theorem 3.7, for given $A$ and $B$, the optimal solution $X^{*}$ that satisfies $||AX^{*}-B||=\min_{X}||AX-B||$ is given by:
@@
$\mathbb{E}(||AA^{+}B-B||^{2})=\mathbb{E}(||(I-AA^{+})(-B)||^{2})=\mathbb{E}(||(I-AA^{+})B||^{2})=\sum_{i=1}^{n}\mathbb{E}(||(I-AA^{+})\mathbf{b}_{i}||^{2})$ (38)
According to Lemma G.3 and Corollary G.4, $I-AA^{+}$ is an orthogonal projection and $\text{rank}(I-AA^{+})=m-\min(m-n)$. With Lemma G.5, we know that for any $i$, $\mathbb{E}(||(I-AA^{+})\mathbf{b}_{i}||^{2})=m-\min(m-n)$. Therefore, we can conclude that:
$\mathbb{E}(||AA^{+}B-B||^{2})=\begin{cases}n(m-n),&\text{if }m\geq n,\\ 0,&\text{if }m<n.\end{cases}$ (39)
which leads to the conclusion in Theorem G.1. ∎

### v2

@@ Appendix I More Insights on Linear Compensation
[…]
@@
We give the proof for Theorem I.1 in the following subsections.
I.1 Preliminaries
To prove Theorem I.1, we adopt the notations in Theorem I.1 and prove a series of lemmas first.
Lemma I.2.
Let $AA^{+}, PP$ is an orthogonal projection and $\text{Im}(P)=\text{Col}(A)$.
@@ I.1 Preliminaries
Lemma I.3.
IPI-P is an orthogonal projection and $\text{Im}(I-P)=\text{Col}(A)^{\perp}$.
@@
First, we prove that IPI-P is an orthogonal projection:
@@
$\langle\mathbf{v},\mathbf{d}\rangle$ $=\langle(I-P)\mathbf{c},A\mathbf{w}\rangle=\langle\mathbf{c},(I-P)A\mathbf{w}\rangle$ (35)
$=\langle\mathbf{c},(A-AA^{+}A)\mathbf{w}\rangle=\langle\mathbf{c},0\rangle=0$
@@
Conversely, suppose $\mathbf{u}\in\text{Col}(A)^{\perp}$. Recall Lemma I.2 that $P$ is an orthogonal projection and $\text{Im}(P)=\text{Col}(A)$, we have $P\mathbf{u}=0$. Therefore,
@@
From Lemma I.3, we can immediately conclude that:
Corollary I.4.
@@
$\text{rank}(I-P)=\dim\text{Im}(I-P)=\dim\text{Col}(A)^{\perp}=m-\text{rank}(A)$. Recall Theorem I.1 that every element from $A$ is sampled from a Gaussian distribution, so that by probability of 1 we have $\text{rank}(A)=\min(m,n)$, which leads to the conclusion of this corollary. ∎
@@
[…]
@@
[…]
I.2 Proof
Finally, we give the proof for Theorem I.1.
@@ I.2 Proof
[…]
@@
[…]
which leads to the conclusion in Theorem I.1. ∎